OpenStudy (anonymous):

roots of lx-3l+lxl-6

OpenStudy (anonymous):

There are 3 or 1 possible negative roots

OpenStudy (anonymous):

saifooooooooooo! you need a nap

OpenStudy (saifoo.khan):

=(

OpenStudy (saifoo.khan):

DANG!

OpenStudy (anonymous):

if x > 3 this is \[x-3+x-6=0\] \[2x=9\] \[x=\frac{9}{2}\] that is one answer

OpenStudy (saifoo.khan):

second is -3/2

OpenStudy (anonymous):

if x <0 this is \[-x+3-x-6=0\] \[-2x=3\] \[x=-\frac{3}{2}\]

OpenStudy (anonymous):

and if 0 < x < 3 this is \[-x+3+x-6=0\] or \[-3=0\] so there is no solution on that intervals

OpenStudy (anonymous):

steps clear or not? you need to break up the absolute value into intervals where it is positive and negative. then get rid of absolute value signs and solve the linear equation